OpenStudy (anonymous):

find the solution set of this inequality. x(x+1) ≤ 20 so i have to take out the 20 first ?

terenzreignz (terenzreignz):

Try turning it into an inequality with 0 on one side.

OpenStudy (anonymous):

solution please! :D

terenzreignz (terenzreignz):

ok, add (-20) to both sides of the inequality, that yields x(x + 1) - 20 ≤ 0 x^2 + x - 20 ≤ 0 (x + 5)(x - 4) ≤ 0 Could you take it from there? :)

OpenStudy (anonymous):

what do you mean by "take out the 20 first"? this is the way I would proceed \[x(x+1)\leq20\] first determine where the expression is equal to zero, these are the"x" values where the statement will change from true to false or visa versa so I want to solve for x in \[x(x+1)=20\] \[x^2+x-20=0\] factoring \[(x+5)(x-4)=0\] so \[x=-5\] or \[x=4\] now test the original statement for intervals of truth\[(x+5)(x-4)\leq 0\] as test numbers I choose \[-6, 0, and, 5\] -6 results in a false statement 0 results in a true statement and 5 results in a false statement so the solution set is the closed interval\[[-5,4]\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

oh yeah i forgot 5&4. geez